OpenStudy (anonymous):

find the product:(5/8)*(3/10)

OpenStudy (anonymous):

okay well when your multiplying factions you simply multiply across so to find the numerator you need to find the product of the two numerators you have (5 * 3) and then do the same thing but with the denominator (8*10) so what is (5*3)

OpenStudy (anonymous):

15

OpenStudy (irishboy123):

\(\large \frac{A}{X} \times \frac{B}{Y} = \frac{A \times B}{X \times Y}\)